OpenStudy (anonymous):

lim of (1 + 7^(n)) / 9^(n) as n ----> oo

OpenStudy (amistre64):

\[\frac{1}{9^n}+(\frac{7}{9})^n\]

OpenStudy (amistre64):

7/9 is a fraction between 0 and 1 so it goes to 0 same with the 1/9 .....

OpenStudy (anonymous):

it goes to zero because its a fraction?

OpenStudy (amistre64):

well, becasue its a fraction whose absolute value is between 0 and 1

OpenStudy (amistre64):

at large n we get: \[\frac{1}{100000000000000}+\frac{puny}{100000000000000}\] which are pretty small

OpenStudy (anonymous):

ahh the way I see it is that the denominator is bigger than the numerator so the fraction will get smaller as n increases. is that correct?

OpenStudy (amistre64):

thats correct

OpenStudy (anonymous):

cool. thanks man. I might be posting another question soon lol

OpenStudy (amistre64):

good luck :)